OpenStudy (anonymous):

Find the solution of the system of equations:x=4+y, 9x=10y=19

OpenStudy (sawpnilsrabon):

First Given 9x=10y=19 9x=19 x=19/9 x=2.11 and 10y=19 y=19/10 y=1.9 Now given again x=4+y x=4+1.9 [y=1.9] x=5.9 and x=4+y 2.11=4+y y=2.11-4[x=2.11] y=-1.89 Ans: x, y= 5.9, 1.89

OpenStudy (anonymous):

i'm confused at this

OpenStudy (sawpnilsrabon):

whats the part confused??